# FEDONet: Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v3

@@ Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations. Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this limitation, we introduce Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-Embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson, Burgers’, Lorenz-63, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation. FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
@@
organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
@@ 1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13] However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
@@
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [14] pioneered the branch-trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [15, 16, 17] extended these ideas with rigorous operator-approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [18] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Related ideas have been explored in the context of learning PDE dynamics in modal or Fourier space [19], where the evolution operator is approximated directly on spectral coefficients rather than physical fields, highlighting the efficacy of frequency-space representations for capturing multiscale dynamics. Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
@@
Beyond architectural and kernel innovations, recent research has expanded operator learning along multiple frontiers. Efforts addressing solution irregularities have focused on handling discontinuities [26], incorporating derivative supervision for enhanced gradient fidelity [27], and enabling learning on irregular meshes through graph-based trunk networks [28]. Adaptive and multi-resolution strategies have been introduced to improve sampling efficiency [29], while hybrid encoder-decoder frameworks leveraging U-Net and wavelet representations further enhance multi-scale expressivity [30, 31]. To embed physical constraints and quantify uncertainty, the paradigm of Physics-Informed Neural Networks (PINNs) [32, 33] has been generalized into Physics-Informed Neural Operators [34, 35] and related variational or pseudo-physics extensions [36, 37]. Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [38, 39, 40, 41]. Transformer based attention mechanisms have also been adapted for operator learning [42, 43, 44, 45, 46] and for robust geometry generalization [47]. Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [48] and its amortized Fourier Neural Operator successor [49], underscoring the maturity and scalability of the operator-learning paradigm.
@@
Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [50], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP. To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [50]. This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters. FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets. A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Eikonal, and Lorenz-63. By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation. Section  2 presents the formulation of the Fourier-Embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity. Section  3 describes the benchmark setup and implementation details across the seven PDE families considered. Sections  4 and  5 summarize the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
@@ 2 Methodology
We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces. Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces:
@@
In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$, comprising function pairs sampled from $\mathcal{G}$. DeepONet is a neural operator architecture well-suited to this task. It decomposes the learning problem into two components: Branch network $B_{\theta}$ that encodes discrete samples of $u$ at sensor points $\{x_{1},\dots,x_{m}\}$, Trunk network $T_{\theta}$ that takes in spatial (or spatiotemporal) coordinates $\zeta=(x,y,z,t)$.
The predicted value at location $\zeta$ is computed via an inner product:
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias - they learn low-frequency components faster than high-frequency ones [51]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier Embeddings into the trunk network input illustrated in Fig. 1. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
@@
where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
The FEDONet prediction then becomes:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
Figure 1: FEDONet: Fourier Embedded Deep operator Network
This transformation can be interpreted as a randomized kernel approximation:
@@
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [52], such embeddings approximate shift-invariant kernels - implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
Beyond expressivity, Fourier embeddings improve the conditioning of the learning process. Under mild assumptions, the embedded features exhibit approximate whitening ( see A ) :
@@
reducing feature correlations and leading to a more isotropic Neural Tangent Kernel (NTK) spectrum. This in turn improves optimization dynamics, stabilizes gradient flow, and accelerates convergence during training.
Let $\mathcal{H}_{\phi}$ denote the function class learnable with Fourier features, and $\mathcal{H}_{\text{vanilla}}$ the class learnable with raw coordinates. Then:
$\mathcal{H}_{\phi}\supset\mathcal{H}_{\text{vanilla}},$ (8)
with the inclusion being strict in problems with rich frequency content (see B). Fourier embeddings thus act as spectral preconditioners that align the input geometry with the task’s spectral demands.
@@ 2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
Let the dataset be denoted by
$\mathcal{D}=\left\{\left(u^{(i)},\left\{\zeta^{(i)}_{j},s^{(i)}_{j}\right\}_{j=1}^{Q}\right)\right\}_{i=1}^{N},$
where $N$ is the number of input functions and $Q$ is the number of query points per function.
The model prediction at location $\zeta$ is given by:
$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)),$
where $b_{k}(u)$ and $t_{k}(\phi(\zeta))$ denote the outputs of the branch and Fourier-embedded trunk networks, respectively.
The training objective is to minimize the empirical risk using the mean squared error (MSE) loss:
@@
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function-query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training. This design choice ensures low overhead while enhancing model expressivity through spectral enrichment. Nonetheless, the architecture can be extended to incorporate learnable embeddings $\phi_{\theta}(\zeta)$, enabling adaptive spectral tuning tailored to problem-specific frequency content.
2.3 Evaluation and Generalization
After training, the FEDONet model is evaluated on unseen input functions $u_{\text{test}}$. To assess model performance, we report both quantitative and qualitative metrics:

### v4

@@ Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations. Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this limitation, we use Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random Fourier features to enrich spatial representation capabilities. The Fourier-Embedded DeepONet (FEDONet) demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the Burgers’, 2D Poisson, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation. To systematically evaluate the effectiveness of the architectures, we perform comparisons across multiple training dataset sizes and input noise levels. FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems. This work demonstrates the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
@@
[aff1]organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
@@ 1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element, and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks [4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel-based Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13]. However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long-standing challenge. From a numerical analysis perspective, this difficulty can often be interpreted as a mismatch between the spectral content of PDE solutions and the coordinate representations supplied to standard neural architectures. In such cases, the network must learn highly oscillatory or multiscale structures solely through nonlinear transformations, which can lead to inefficient optimization and poor sample efficiency.
@@
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [14] pioneered the branch-trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [15, 16, 17] extended these ideas with rigorous operator-approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [18] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Related ideas have been explored in the context of learning PDE dynamics in modal or Fourier space [19], where the evolution operator is approximated directly on spectral coefficients rather than physical fields, highlighting the efficacy of frequency-space representations for capturing multiscale dynamics. Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability. Despite these advances, a key challenge remains the development of operator learning architectures that balance spectral accuracy, computational efficiency, and architectural simplicity, particularly for integration into large-scale scientific workflows such as turbulence modeling and data assimilation in fluid mechanics.
@@
Recent research has expanded operator learning along multiple frontiers. Efforts addressing solution irregularities have focused on handling discontinuities [26], incorporating derivative supervision for enhanced gradient fidelity [27], and enabling learning on irregular meshes through graph-based trunk networks [28]. Adaptive and multi-resolution strategies have been introduced to improve sampling efficiency [29], while hybrid encoder-decoder frameworks leveraging U-Net and wavelet representations further enhance multi-scale expressivity [30, 31]. To embed physical constraints and quantify uncertainty, the paradigm of Physics-Informed Neural Networks (PINNs) [32, 33] has been generalized into Physics-Informed Neural Operators [34, 35] and related variational or pseudo-physics extensions [36, 37]. Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [38, 39, 40, 41]. Transformer-based attention mechanisms have also been adapted for operator learning [42, 43, 44, 45, 46] and for robust geometry generalization [47]. Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [48] and its amortized Fourier Neural Operator successor [49], underscoring the maturity and scalability of the operator-learning paradigm. Recent efforts have also begun to explore the application of data-driven operator learning and spectral modeling approaches to more complex fluid dynamics systems, highlighting the importance of robustness, spectral fidelity, and generalization in realistic regimes [50, 51].
@@
Despite recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net-based trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [52], yet their potential within operator-learning frameworks remains largely unexplored, particularly in settings where the trunk is typically implemented as a simple fully connected MLP. From this viewpoint, enriching the coordinate representation with sinusoidal embeddings can be interpreted as a form of spectral lifting that aligns the input representation with the frequency content of typical PDE solutions. Recent studies have also explored incorporating Fourier-based representations within DeepONet frameworks for surrogate modeling in specific application domains [53]. In contrast to such approaches that rely on predefined Fourier bases tailored to particular physical settings, the present work introduces randomized Fourier feature embeddings applied directly to the trunk coordinates. This strategy acts as a lightweight and general-purpose coordinate transformation that enriches the spectral representation capacity of the trunk network while preserving the standard DeepONet formulation. The Fourier-Embedded DeepONet, FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, highlighting its role as a spectral preconditioning mechanism that improves optimization behavior and training stability. This analysis demonstrates that the embedding strictly enlarges the approximation class of standard DeepONets. We conduct a comprehensive benchmark spanning steady 2D Poisson, 1D Viscous Burgers, 1D Allen-Cahn, 1D Kuramoto-Sivashinsky, and 2D Eikonal equation. In addition to accuracy comparisons, we perform systematic analyses examining the dependence of model performance on training dataset size and input noise levels, providing further insight into the robustness and data efficiency of the architectures.
By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation. The primary contributions of this work are summarized as follows. First, we introduce a Fourier-Embedded Deep Operator Network (FEDONet) that enriches trunk coordinate representations through randomized Fourier feature embeddings, providing a lightweight spectral lifting mechanism for DeepONet architectures. Second, we provide theoretical insight into the effect of this embedding through the operator neural tangent kernel, showing that the resulting architecture expands the effective hypothesis space of standard DeepONets while improving optimization conditioning. Third, we perform a systematic empirical evaluation across multiple canonical PDE benchmarks and analyze robustness with respect to dataset size and input noise levels. Together, these results demonstrate that spectral coordinate embeddings provide a simple yet effective mechanism for improving operator learning performance without increasing architectural complexity.
@@ 2 Methodology
We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces. Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces,
@@
In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$ comprising function pairs sampled from $\mathcal{G}$. DeepONet is a neural operator architecture well-suited to this task. It decomposes the learning problem into two components,
  • Branch network $B_{\theta}$ that encodes discrete samples of the input function $u$ evaluated at sensor locations $\{x_{1},\dots,x_{m}\}$.
  • Trunk network $T_{\theta}$ that receives spatial or spatiotemporal coordinates $\zeta\in\mathcal{Y}$ (e.g., $\zeta=(x,y,z,t)$).
The predicted value of the operator at location $\zeta$ is computed through an inner product between the branch and trunk outputs,
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
Standard multilayer perceptrons (MLPs) exhibit a well-known spectral bias, whereby low-frequency components are learned significantly faster than high-frequency components [54]. This phenomenon becomes particularly problematic in operator learning tasks involving sharp gradients, multiscale dynamics, or oscillatory solution fields. From a numerical perspective, this limitation can be viewed as a conditioning issue arising from an inadequate coordinate representation. When raw spatial coordinates are used as inputs, the network must internally construct oscillatory features through deep nonlinear transformations, which can slow optimization and reduce sample efficiency.
To address this limitation, we use fixed Fourier Embeddings applied to the trunk network input, illustrated in Fig. 1. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
@@
where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk network, we provide the transformed features $\phi(\zeta)$, thereby enriching the spectral representation of the coordinate inputs. Unlike Fourier-basis DeepONet [53] approaches that construct the input function space using predefined Fourier modes, the present formulation does not impose any explicit spectral structure on the input functions themselves. Instead, Fourier embeddings act purely as a coordinate transformation that enhances the representational capacity of the trunk network while preserving the standard DeepONet architecture.
The FEDONet prediction therefore becomes
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\,t_{k}(\phi(\zeta)).$ (5)
Figure 1: FEDONet: Fourier-Embedded Deep Operator Network.
This transformation can be interpreted as a randomized kernel approximation
@@
which lifts the coordinate $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [55], such embeddings approximate shift-invariant kernels and therefore implicitly kernelize the trunk network. This spectral lifting enables the trunk network to represent oscillatory and multiscale spatial patterns using shallow architectures, improving the learning efficiency for PDE operators with rich frequency content.
Beyond representational expressivity, Fourier embeddings also improve the conditioning of the learning problem. Under mild assumptions, the embedded features exhibit approximate whitening (see A):
@@
which reduces feature correlations and produces a more isotropic Neural Tangent Kernel (NTK) spectrum. This improved conditioning leads to more stable gradient propagation and faster convergence during training, effectively acting as a spectral preconditioning mechanism for operator learning.
@@ 2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions. Each dataset element consists of an input function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
The dataset can therefore be written as
$\mathcal{D}=\left\{\left(u^{(i)},\{\zeta^{(i)}_{j},s^{(i)}_{j}\}_{j=1}^{Q}\right)\right\}_{i=1}^{N}.$
The model prediction at location $\zeta$ is
$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\,t_{k}(\phi(\zeta)).$
The training objective minimizes the empirical risk using the mean squared error loss:
@@
Model parameters are optimized using mini-batch optimization using Adam. Each mini-batch consists of randomly sampled function-coordinate pairs $(u,\zeta)$, which enables scalable training and improves generalization across diverse operator inputs. The Fourier embedding layer $\phi(\zeta)$ remains fixed and non-trainable throughout training, ensuring minimal computational overhead while enriching the coordinate representation. Although the present work focuses on fixed random embeddings, the framework can naturally be extended to learnable embeddings $\phi_{\theta}(\zeta)$ that adapt the spectral representation to problem-specific frequency content.
2.3 Evaluation
After training, DeepONet and FEDONet are evaluated on unseen input functions $u_{\text{test}}$ sampled from the same distribution as the training set. The evaluation is designed to probe not only pointwise accuracy but also the model’s ability to capture multiscale structures, remain robust under varying noise levels, and generalize effectively in limited-data regimes. To assess model performance, we report both quantitative and qualitative metrics: